OpenStudy (anonymous):

6/5x - 1/x+1=1/3x^2+3x HELP!!

OpenStudy (anonymous):

You could start by solving the right side using the quadratic equation...That might be tricky though.

OpenStudy (anonymous):

im having trouble solving it and dont know what to do

OpenStudy (mathstudent55):

Can you use the equation editor or the drawing tool? It's hard to understand what you mean.

OpenStudy (mathstudent55):

You may have meant this, but this is definitely not what you wrote above. \(\dfrac{6}{5x} - \dfrac{1}{x + 1} = \dfrac{1}{3x^2 + 3x} \)

OpenStudy (anonymous):